OpenStudy (anonymous):

Given the function f(x) = log5(x + 1), find the value of f−1(2). @Hitaro9

OpenStudy (hitaro9):

Oh. Is that an inverse function?

OpenStudy (isaiah.feynman):

http://openstudy.com/users/isaiah.feynman#/updates/52b13d8ee4b03e81ff93c5b9 See my solution there.

OpenStudy (anonymous):

@Hitaro9 yes I think so

OpenStudy (isaiah.feynman):

This question is no different from yesterday's.

OpenStudy (isaiah.feynman):

Procedure to find the inverse of a function. (This what I did) Change f(x) to y switch x and y solve for y the new function is the inverse. Follow them carefully Sara :)

OpenStudy (anonymous):

f^-1= 5(2)-1?

OpenStudy (anonymous):

@Isaiah.Feynman like that I dont know what to do after that

OpenStudy (isaiah.feynman):

Actually you are done.

OpenStudy (anonymous):

f−1(2) = 3 f−1(2) = 11 f−1(2) = 18 f−1(2) = 24 these are the options and its like none of these

OpenStudy (isaiah.feynman):

Its 5^2 not 5*2

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh (thats the sound of my epiphany)

OpenStudy (anonymous):

so its the last one

OpenStudy (isaiah.feynman):

Yep.